OpenStudy (anonymous):

9q-14+3(q-8)=7 step by step find q

OpenStudy (anonymous):

9q-14+3(q-8)=7 9q-14+3q-24=7 9q+3q=7+14+24 12q=45 12q/12=45/12 q=3.75